OpenStudy (anonymous):

Given f(x) = 6x and g(x) = 2x +1 find f(2) + g(6)

OpenStudy (fibonaccichick666):

do you know how to find f(2) by itself?

OpenStudy (anonymous):

not really.

OpenStudy (fibonaccichick666):

ok, it's not bad. It just means that wherever you see an x in f(x) you substitute in a 2 instead. For example h(x)=5x+2 find h(3) h(3)=5(3)+2=15+2=17

OpenStudy (fibonaccichick666):

will you try finding f(2)?

OpenStudy (anonymous):

so it would be f(2)=6(2)+G(2)=2(2)+1?

OpenStudy (fibonaccichick666):

wait, just find f(2) to start

OpenStudy (fibonaccichick666):

what is your f(x)?

OpenStudy (anonymous):

WAIT NOW IM CONFUSED

OpenStudy (anonymous):

Ignore the caps my bad

OpenStudy (fibonaccichick666):

so, what is f(x) in your question?

OpenStudy (fibonaccichick666):

ie what is it equal to?